# Exploring the evolution of red and blue galaxies in different cosmic web environments using IllustrisTNG simulation

### v1

@@ Abstract
We analyze the evolution of red and blue galaxies in different cosmic web environments from redshift $z=3$ to $z=0$ using the IllustrisTNG simulation. We use Otsu’s method to classify the red or blue galaxies at each redshift and determine their geometric environments from the eigenvalues of the deformation tensor. Our analysis shows that initially, blue galaxies are more common in clusters followed by filaments, sheets and voids. However, this trend reverses at lower redshifts, with red fractions rising earlier in denser environments. At $z<1$, most massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) are quenched across all environments. In contrast, low-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}})<10.5$) are more influenced by their environment, with clusters hosting the highest red galaxy fractions at low redshifts. We observe a slower mass growth for low-mass galaxies in clusters at $z<1$. Filaments show relative red fractions (RRF) comparable to clusters at low masses, but host nearly $60\%$ of low-mass blue galaxies, representing a diverse galaxy population. It implies that less intense environmental quenching in filaments allows galaxies to experience a broader range of evolutionary stages. Despite being the densest environment, clusters display the highest relative blue fraction (RBF) for high-mass galaxies, likely due to interactions or mergers that can temporarily rejuvenate star formation in some of them. The $(u-r)$ colour distribution transitions from unimodal to bimodal by redshift $z=2$ across all environments. At $z<1$, clusters exhibit the highest median colour, with stellar mass being the primary driver of colour evolution in massive galaxies. The suppression of star formation rate (SFR) and specific SFR (sSFR) is also most pronounced in clusters during this period. Our study suggests that stellar mass governs quenching in high-mass galaxies, while a complex interplay of mass and environment shapes the evolution of low-mass galaxies.
@@ 2 Data
IllustrisTNG [^1] https://www.tng-project.org/ [106, 107, 102, 104, 105, 103] is a suite of cosmological gravo-magnetohydrodynamical simulations built on the moving-mesh code AREPO [116, 117]. It serves as an enhanced version of the original Illustris project [118, 119, 120]. The TNG suite consists of three cosmological volumes with side lengths of approximately 50, 100, and 300 Mpc, named TNG50, TNG100, and TNG300, respectively. Each volume offers multiple resolution levels, such as TNG100-1, TNG100-2, TNG100-3, and similarly for TNG300, whereas TNG50 has four distinct resolution outputs. All simulations begin at a redshift of $z=127$, using cosmological parameters from the Planck 2015 results [121]: $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$. For our analysis, we utilize the TNG300-1 run, which offers the largest volume and highest resolution in the TNG series. The specifics of TNG300-1 are detailed in the  Table 1, and our study covers redshifts from $z=3$ to the present ($z=0$).
@@
At each redshift, we extract various galaxy properties from the publicly available group catalog, which provides details on both halos and subhalos. The FoF algorithm is used to identify halos, while the subhalos (or galaxies) are identified using the SUBFIND algorithm [122, 123]. In this work, we select only galaxies with a non-zero SubhaloFlag, as a value of zero indicates a subhalo of non-cosmological origin. Additionally, we impose a stellar mass criterion, choosing galaxies in the range $129<\log_{10}(\frac{M_{\star}}{M_{\odot}})<$ 12. Stellar mass is defined as the total mass of star particles within twice the stellar half-mass radius $(r_{stars},1/2)$ [105]. For $(u-r)$ colour, we use the supplementary catalog provided by  [102]. The star formation rates are obtained from the SubhaloSFRinRad field in the group catalog. In IllustrisTNG, star formation is modeled following  [124]. For our analysis, we use the instantaneous star formation rate, which is the sum of the star formation rates of all star-forming gas cells within twice the stellar half-mass radius of the galaxy.
Figure 1: This figure shows the distribution of galaxies in the $(u-r)$ colour-stellar mass plane across each redshift, with the dividing line (determined by Otsu’s method) distinguishing the red and blue populations.
@@ 3.1 Identifying red and blue Galaxies with Otsu’s method
Otsu’s thresholding technique [115], originally designed for image segmentation, separates data into two groups by minimizing the intra-class variance (variance within each group) and maximizing the inter-class variance (variance between groups). This method iterates through all possible thresholds, selecting the one that provides the optimal separation between the two classes. This technique is particularly useful when the data distribution is bimodal, as is the case with the $(u-r)$ colour distribution of galaxies, where we observe two peaks corresponding to the blue cloud and the red sequence. Recently, [114] proposed a method for classifying galaxies into two groups: the “blue cloud” (younger, star-forming galaxies) and the “red sequence” (older, passive galaxies) based on the Otsu’s technique.
@@
    Histogram Calculation: We first calculate the histogram of the $(u-r)$ colour for all galaxies in our sample, using a specific number of bins $M$. The histogram is then normalized by the total number of galaxies $N=\sum_{i=1}^{M}n_{i}$ where $n_{i}$ corresponds to the number of galaxies in the $i^{th}$ colour bin. This ensures that the sum of the probabilities for all bins equals to one i.e $\sum_{i=1}^{M}p_{i}=1$ where $p_{i}=\frac{n_{i}}{N}$.
@@
    Class Probabilities: For a given threshold, the galaxies are divided into two groups: blue cloud (BC) and red sequence (RS). The probability of each class occurring is calculated based on the proportion of galaxies in each group relative to the total sample. If the threshold is at the k-th bin, the blue cloud is represented by the first $k$ bins, while the red sequence is represented by the bins from $k+1$ to the end. The class probabilities are given by,
@@
    Here, $x_{i}$ represents the $(u-r)$ colour corresponding to the $i^{th}$ bin. The mean up to the $k^{th}$ bin is given by $\mu_{k}=\sum_{i=1}^{k}x_{i}p_{i}$, while $\mu_{T}=\sum_{i=1}^{M}x_{i}p_{i}$ denotes the mean of the entire distribution. It is important to note that $P_{BC}+P_{RS}=1$ and that $\mu_{T}=P_{BC}\,\mu_{BC}+P_{RS}\,\mu_{RS}$ for every threshold.
@@
    Intra-class variance and Inter-class variance: The intra-class (within-class) variance ($\sigma_{wc}^{2}$) and the inter-class (between-class) variance ($\sigma_{bc}^{2}$) are calculated as,
@@
    Threshold Selection: The optimal threshold is the one that minimizes the intra-class variance $\sigma_{wc}^{2}$, which is equivalent to maximizing the inter-class variance $\sigma_{bc}^{2}$. In practice, this threshold can be found by iterating through the possible values and selecting the one that satisfies these conditions. This yields a threshold that is not affected by binning choices, making it a reliable approach for classifying red and blue galaxies [114]. This method offers a robust and parameter-free way to classify galaxies into the blue cloud and red sequence based on their $(u-r)$ colour and stellar mass.
@@
All $9<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)\leq 10.5$ $1210.5<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)<$ 12
@@ 00 $253140$ $130751$ $122389$ $225075$ $105417$ $119658$ $28065$ $25334$ $2731$
$2$ $157394$ $66878$ $90516$ $147377$ $61740$ $85637$ $10017$ $5138$ $4879$
$3$ $84438$ $46230$ $38208$ $80992$ $44559$ $36433$ $3446$ $1671$ $1775$
Table 2: This table shows the total number of galaxies with nonzero SubhaloFlag and stellar masses within $129<\log_{10}(\frac{M_{\star}}{M_{\odot}})<$ 12 in IllustrisTNG simulation for each snapshot from redshift 00 to $3$. It also lists the number of red and blue galaxies identified with the Otsu’s method at each redshift in different mass ranges.
There are distinct relationships between colour and stellar mass or absolute magnitude. Therefore, a single colour threshold cannot be justified for galaxies of varying masses and luminosities. We divide the entire sample into several independent stellar mass bins and apply this technique to each of these bins separately. This provides us a dividing line between the two populations in the colour-stellar mass plane at each redshift (Figure 1). We also tabulate the number of red and blue galaxies in two different mass ranges between redshift $3$ to 00 in Table 2.
Figure 2: This figure displays the distribution of galaxies in various cosmic web environments at $z=0$, from a 20 Mpc thick slice of the IllustrisTNG simulation.
@@ 3.2 Identifying different morphological environments of the cosmic web
The deformation tensor $T_{ij}$ is derived from the Hessian matrix of the gravitational potential field $\Phi$, defined as:
@@
Here, $\delta=\frac{\rho-\bar{\rho}}{\bar{\rho}}$ is the density contrast, with $\rho$ representing the local density and $\bar{\rho}$ the average density. To calculate the potential, we apply the Cloud-In-Cell (CIC) scheme to construct a discrete density contrast field on a grid. This field is smoothed using an isotropic Gaussian filter with a width of 4 Mpc.

### v2

@@ Abstract
We analyze the evolution of red and blue galaxies in different cosmic web environments from redshift $z=3$ to $z=0$ using the IllustrisTNG simulation. We use Otsu’s method to classify the red or blue galaxies at each redshift and determine their geometric environments from the eigenvalues of the deformation tensor. Our analysis shows that initially, blue galaxies are more common in clusters followed by filaments, sheets and voids. However, this trend reverses at lower redshifts, with red fractions rising earlier in denser environments. At $z<1$, most massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) are quenched across all environments. In contrast, low-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}})<10.5$) are more influenced by their environment, with clusters hosting the highest red galaxy fractions at low redshifts. We observe a slower mass growth for low-mass galaxies in clusters at $z<1$. Filaments show relative red fractions (RRF) comparable to clusters at low masses, but host nearly $60\%$ of low-mass blue galaxies, representing a diverse galaxy population. It implies that less intense environmental quenching in filaments allows galaxies to experience a broader range of evolutionary stages. Despite being the densest environment, clusters display the highest relative blue fraction (RBF) for high-mass galaxies, likely due to interactions or mergers that can temporarily rejuvenate star formation in some of them. The (ur)(u-r) colour distribution transitions from unimodal to bimodal by redshift $z=2$ across all environments. At $z<1$, clusters exhibit the highest median colour, with stellar mass being the primary driver of colour evolution in massive galaxies. The suppression of star formation rate (SFR) and specific SFR (sSFR) is also most pronounced in clusters during this period. Our study suggests that stellar mass governs quenching in high-mass galaxies, while a complex interplay of mass and environment shapes the evolution of low-mass galaxies.
@@ 2 Data
IllustrisTNG [^1] https://www.tng-project.org/ [102, 103, 98, 100, 101, 99] is a suite of cosmological gravo-magnetohydrodynamical simulations built on the moving-mesh code AREPO [112, 113]. It serves as an enhanced version of the original Illustris project [114, 115, 116]. The TNG suite consists of three cosmological volumes with side lengths of approximately 50, 100, and 300 Mpc, named TNG50, TNG100, and TNG300, respectively. Each volume offers multiple resolution levels, such as TNG100-1, TNG100-2, TNG100-3, and similarly for TNG300, whereas TNG50 has four distinct resolution outputs. All simulations begin at a redshift of $z=127$, using cosmological parameters from the Planck 2015 results [117]: $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$. For our analysis, we utilize the TNG300-1 run, which offers the highest resolution run among the largest available volume in the TNG series. The specifics of TNG300-1 are detailed in the  Table 1, and our study covers redshifts from $z=3$ to the present ($z=0$).
@@
At each redshift, we extract various galaxy properties from the publicly available group catalog, which provides details on both halos and subhalos. The FoF algorithm is used to identify halos, while the subhalos (or galaxies) are identified using the SUBFIND algorithm [118, 119]. In this work, we select only galaxies with a non-zero SubhaloFlag, as a value of zero indicates a subhalo of non-cosmological origin. Additionally, we impose a stellar mass criterion, choosing galaxies in the range $9<\log_{10}(\frac{M_{\star}}{M_{\odot}})<12$. Stellar mass is defined as the total mass of star particles within twice the stellar half-mass radius $(r_{stars},1/2)$ [101]. For (ur)(u-r) colour, we use the supplementary catalog provided by [98], which provides apparent magnitudes for different snapshots and dust-corrected absolute magnitudes only at $z=0$. At $z=0$, we use the dust-corrected absolute magnitudes to compute colours. For all other redshifts ($z>0$), we compute colours using the apparent magnitudes, as dust-corrected values are not available. Since colour is determined by the difference between magnitudes in two bands, and the distance modulus is the same for both bands of a given galaxy, the colour remains a relative quantity. Although the absence of dust correction at $z>0$ may introduce systematic offsets, our classification method is designed to be robust to such effects. Specifically, we apply Otsu’s thresholding method [111, 110] to the (ur)(u-r) colour distribution at each redshift and in different stellar mass bins. This data-driven method identifies the optimal colour threshold based on the bimodal distribution of galaxy colours, allowing for a consistent and redshift-dependent separation of red and blue populations without relying on fixed colour cuts or rest-frame corrections. The bimodal distribution is more prominent in (ur)(u-r) than other colours, such as (gr)(g-r). The u-band flux is strongly affected by young, hot stars, while r-band flux traces older stars. This makes (ur)(u-r) colour a better indicator of recent star formation activity, whereas (gr)(g-r) is less sensitive to changes in young stellar populations. We obtain the star formation rates from the SubhaloSFRinRad field in the group catalog. In IllustrisTNG, star formation is modeled following [120]. For our analysis, we use the instantaneous star formation rate, which is the sum of the star formation rates of all star-forming gas cells within twice the stellar half-mass radius of the galaxy.
Figure 1: This figure shows the distribution of galaxies in the (ur)(u-r) colour-stellar mass plane across each redshift, with the dividing line (determined by Otsu’s method) distinguishing the red and blue populations. In each subplot, we also show $10$ uniformly spaced contours lying between the minimum and maximum density of galaxies in the respective colour-stellar mass plane.
@@ 3 Method
The (ur(u-r) colour distribution is bimodal at low redshift, with peaks corresponding to the blue cloud and the red sequence. Traditionally a fixed (ur)(u-r) colour cut is used in many earlier studies to separate the red and blue galaxies. But a fixed colour threshold assumes the same division between red and blue galaxies regardless of redshift or stellar mass, which does not reflect galaxy evolution. The precise location of the peaks of the two populations evolves with redshift and stellar mass. Thus, a static threshold could misclassify galaxies, particularly at higher redshifts where the colour distribution is less bimodal. Otsu’s method [111, 110] is better than a fixed colour cut because it provides adaptive, optimal separation of red and blue galaxies while accounting for variations across different stellar mass bins and redshifts.
@@ 3.1 Identifying red and blue Galaxies with Otsu’s method
Otsu’s thresholding technique [111], originally designed for image segmentation, separates data into two groups by minimizing the intra-class variance (variance within each group) and maximizing the inter-class variance (variance between groups). This method iterates through all possible thresholds, selecting the one that provides the optimal separation between the two classes. This technique is particularly useful when the data distribution is bimodal, as is the case with the (ur)(u-r) colour distribution of galaxies, where we observe two peaks corresponding to the blue cloud and the red sequence. Recently, [110] proposed a method for classifying galaxies into two groups: the “blue cloud” (younger, star-forming galaxies) and the “red sequence” (older, passive galaxies) based on the Otsu’s technique.
@@
    Histogram Calculation: We first calculate the histogram of the (ur)(u-r) colour for all galaxies in our sample, using a specific number of bins $M$. The histogram is then normalized by the total number of galaxies $N=\sum_{i=1}^{M}n_{i}$ where $n_{i}$ corresponds to the number of galaxies in the $i^{th}$ colour bin. This ensures that the total probability across all bins sums to one i.e. $\sum_{i=1}^{M}p_{i}=1$ where $p_{i}=\frac{n_{i}}{N}$.
@@
    Class Probabilities: A threshold ($k$-th bin) is chosen, dividing galaxies into blue cloud (bins $1$ to $k$) and red sequence (bins $k+1$ to $M$). We iterate through all possible thresholds (i.e. $k$ assumes all values between $1$ to M1M-1). The probability of a galaxy belonging to each class is computed. The class probabilities of the blue cloud (BC) and the red sequence (RS) are respectively given by,
@@
    Here, $x_{i}$ represents the (ur)(u-r) colour corresponding to the $i^{th}$ bin. The mean up to the $k^{th}$ bin is given by $\mu_{k}=\sum_{i=1}^{k}x_{i}p_{i}$, while $\mu_{T}=\sum_{i=1}^{M}x_{i}p_{i}$ denotes the mean of the entire distribution. It is important to note that $P_{BC}+P_{RS}=1$ and that $\mu_{T}=P_{BC}\,\mu_{BC}+P_{RS}\,\mu_{RS}$ for every threshold.
@@
    Intra-class variance and Inter-class variance: The total variance is divided into intra-class (within-class) variance and inter-class (between-class) variance. The intra-class variance ($\sigma_{wc}^{2}$) and the inter-class variance ($\sigma_{bc}^{2}$) are calculated as,
@@
    Threshold Selection: The optimal threshold is the one that minimizes the intra-class variance $\sigma_{wc}^{2}$, which is equivalent to maximizing the inter-class variance $\sigma_{bc}^{2}$. We iterate through all possible thresholds, selecting the one that minimizes the intra-class variance while maximizing the inter-class variance. This threshold is not affected by binning choices, making it a reliable approach for classifying red and blue galaxies [110].
@@
All $9<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)\leq 10.5$ $10.5<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)<12$
@@ 11 $231933$ $73719$ $158214$ $210682$ $60488$ $150194$ $21251$ $13231$ $8020$
22 $157394$ $66878$ $90516$ $147377$ $61740$ $85637$ $10017$ $5138$ $4879$
33 $84438$ $46230$ $38208$ $80992$ $44559$ $36433$ $3446$ $1671$ $1775$
Table 2: This table shows the total number of galaxies with nonzero SubhaloFlag and stellar masses within $9<\log_{10}(\frac{M_{\star}}{M_{\odot}})<12$ in IllustrisTNG simulation for each snapshot from redshift 0 to $3$. It also lists the number of red and blue galaxies identified with the Otsu’s method at each redshift in different mass ranges.
There are distinct relationships between colour and stellar mass or absolute magnitude. Therefore, a single colour threshold cannot be justified for galaxies of varying masses and luminosities. We divide the entire sample into several independent stellar mass bins and apply this technique to each of these bins separately, allowing the classification to adapt to how colour varies with mass. This provides us a dividing line between the two populations in the colour-stellar mass plane at each redshift (Figure 1). Otsu’s method determines the optimal division between red and blue population at each redshift by minimizing the intra-class variance and maximizing the inter-class variance, ensuring the separation is statistically robust.
We tabulate the number of red and blue galaxies in two different mass ranges between redshift $3$ to 0 in Table 2.
It may be noted that at higher redshifts, particularly at $z=3$, the (ur)(u-r) colour distribution becomes less clearly bimodal (bottom right panel of Figure 11). One-dimensional histograms at this epoch show a largely unimodal shape, making the identification of distinct red and blue populations less straightforward. However, when viewed in the stellar mass-colour plane, a subtle population of massive galaxies with relatively higher (ur)(u-r) colours emerges, suggesting that a weak form of bimodality may already be developing (bottom right panel of Figure 1). To maintain methodological consistency across redshifts, we continue to apply Otsu’s thresholding method at $z=3$ as a data-driven way to divide the galaxy population into two relative colour classes. At this redshift, the resulting groups are better interpreted as “less blue” and “more blue” galaxies, rather than physically distinct red and blue populations. This functional classification still provides meaningful insights into environmental trends at early times. This approach allows for a uniform comparative analysis across cosmic time, while acknowledging the evolving nature of galaxy colour distributions.
Figure 2: This figure displays the distribution of galaxies in various cosmic web environments at $z=0$, from a 20 Mpc thick slice of the IllustrisTNG simulation. The cosmic web classification is based on the eigenvalues of the deformation tensor, estimated from the density field smoothed with a Gaussian filter of width 4 Mpc.
@@ 3.2 Identifying different morphological environments of the cosmic web
The deformation tensor $T_{ij}$ is derived from the Hessian matrix of the gravitational potential field $\Phi$, defined as
@@
Here, $\delta=\frac{\rho-\bar{\rho}}{\bar{\rho}}$ is the density contrast, with $\rho$ representing the local density and $\bar{\rho}$ the average density. To calculate the potential, we apply the Cloud-In-Cell (CIC) scheme to construct a discrete density contrast field on a grid. This field is smoothed using an isotropic Gaussian filter with a width of 4 Mpc. The mean intergalactic separation of our samples varies between $4.79$ Mpc (at $z=0$) to $6.91$ Mpc (at $z=3$). Further, we have calculated the potential using a discretized density field with a grid spacing of $\sim 1.18$ Mpc. The smoothing scale should be larger than the grid spacing to ensure meaningful results. We are primarily concerned about resolving large-scale structures of the cosmic web. So, we have chosen a filter width of 4 Mpc for our present analysis. We also study the effects of varying the filter width in subsection A.2.
@@
It may be worthwhile to mention that the density field used in this study is based on the spatial distribution of galaxies, rather than the underlying dark matter field. While the large-scale structure is fundamentally shaped by the dark matter distribution, and galaxies are biased tracers of that field, our choice reflects the practical constraints of observational studies. Since dark matter is not directly observable, galaxy-based density fields provide a more realistic and consistent basis for comparison with observational data. Our focus on galaxy properties motivates the use of a density measure derived from the same population, facilitating a more direct link between environment and observable galaxy characteristics. We utilize a subset of subhalos in each redshift applying the criteria discussed in section 2. The total number of subhalos in each redshift are presented in the first sub-column (“Total”) of the first column (“All”) of Table 2.
Redshift Number of Galaxies
Total Void Sheet Filament Cluster
0 253140 1742 40510 128945 81943
0.1 253212 1746 40213 129829 81424
0.2 252546 1728 39965 130091 80762
0.3 251329 1729 39453 129785 80362
0.4 249659 1749 38779 129826 79305
0.5 247760 1725 38092 128924 79019
0.7 242859 1668 36612 127139 77440
1 231933 1580 34128 121051 75174
1.5 199307 1187 27280 103015 67825
2 157394 850 19497 80358 56689
3 84438 224 7860 41229 35125
Table 3: This table represents the number of galaxies identified in the different regions of the cosmic web at different redshifts.
[…]
Figure 4: This figure shows the distribution of galaxy local densities across different cosmic web environments at each redshift. Here, local density refers to the density estimated using the CIC interpolation scheme applied to the galaxy distribution.
We list the number of galaxies present in each type of cosmic web environments at different redshifts in Table 3. We also show the evolution of the relative fraction of galaxies in each type of cosmic web environments for the entire sample and the two distinct mass bins in Figure 3. The results are obtained for a smoothing length of 4 Mpc. The left panel of Figure 3 shows that filaments generally have the highest galaxy abundance, followed by clusters, sheets, and voids across the entire redshift range. The proportion of galaxies in filaments and sheets increases with decreasing redshift, while the proportion in clusters decreases. These trends are also evident in the lower mass bin, as depicted in the middle panel of Figure 3. In the high mass bin, the proportion of galaxies in filaments and clusters evolves differently, as shown in the right panel of Figure 3. There is a crossover around $z\sim 1$, after which filaments have a higher proportion of galaxies compared to clusters. Prior to $z=1$, clusters dominate in terms of galaxy proportion.
An earlier analysis [78] using the Millennium simulation [123] find that in the present Universe, the mass fraction in filaments and sheets are $\sim 50\%$ and $\sim 24\%$ respectively. Another study [124] with EAGLE simulation [125] find that $\sim 62-66\%$ galaxies with stellar mass larger than 5 times $10^{8}\,M_{\odot}/h$ are in filaments at $z=0$. Despite differences in the method, sample and simulation, it is interesting to note that our analysis indicates $\sim 52\%$ galaxies are in filaments and $\sim 18\%$ galaxies are in sheets at $z=0$.
We also show the probability distribution function (PDF) of the local density across various cosmic web environments at different redshifts in different panels of Figure 4. The peaks of the distributions for different cosmic web environments occur at different densities. The peak for nodes appears at the highest density, followed in order by filaments, sheets, and voids. At each redshift, the low density regions are dominated by voids followed by sheets and filaments. The peak for the nodes exhibit a gradual shift towards higher densities with decreasing redshift.